OpenStudy (anonymous):

Evaluate the limits that exist 1. Lim as x approaches 0 of sin^2 3x/ 5x^2 2. and sin x/2x^2 -x

OpenStudy (anonymous):

\[\lim_{x\to 0}\frac{\sin^23x}{5x^2}\] You're probably familiar with the limit \[\lim_{x\to 0}\frac{\sin ax}{ax} = 1, \text{ right? Here, $a$ is a non-zero constant.}\]

OpenStudy (anonymous):

use l'hopital twice if you are allowed to use it

OpenStudy (anonymous):

I have a similar question that my professor gave me \[\lim_{x \rightarrow 0} \frac{ \sin^3 t }{ (2t)^3 } \] and the answer is 1/8 I believe its the same at that question so i don't think 1 is the answer

OpenStudy (anonymous):

well i may be wrong i don't know

OpenStudy (anonymous):

I'm not saying 1 is the answer. However, you could use the result from that known special limit to evaluate the one (or two) at hand. Specifically, I mean \[\large\lim_{x\to 0} \frac{\sin^23x}{5x^2} =\frac{1}{5}\left(\lim_{x\to 0} \frac{\sin3x}{x} \frac{\sin3x}{x}\right) \] Now, using the known limit, how can you rewrite the right hand side to get something easy to evaluate?

OpenStudy (anonymous):

Ohhhh I understand. Thank you so much.

OpenStudy (anonymous):

Do you also need help with the second limit? It's a bit hard to interpret. Do you mean \[\lim{x\to 0}\frac{\sin x}{2x^2-x}?\]

OpenStudy (anonymous):

And you're welcome :)

OpenStudy (anonymous):

Yes thats what I mean and yes help would be much appreciated!

OpenStudy (anonymous):

Similar to the first one, I suggest factoring out \[\frac{\sin x}{x}\] Then use what you know about limits (namely, their properties) to figure it all out. That's usually the strategy for these kinds of limits, unless you know L'Hopital's rule that @satellite73 mentioned.

OpenStudy (anonymous):

I don't know the L'Hospital rule, but I will definitely google it. Thanks a lot for your help.

OpenStudy (anonymous):

I don't know the L'Hospital rule, but I will definitely google it. Thanks a lot for your help.